OpenStudy (anonymous):

The length of a rectangle is 3 times its width. If the length were increased by 2 inches and the width by 1 inch, the new perimeter would be 62 inches. What is the width and length of the rectangle? {Width in., Length in.}

OpenStudy (anonymous):

The setup equation should be: 2(3x+2)+2(x+1)=62

OpenStudy (anonymous):

If we solve that, the original width equals 7 while the length equals 21.

OpenStudy (anonymous):

Does that help? :)

OpenStudy (ajprincess):

Let the width of the rectangle be w and the length be l. l=3w (it is given that length is three times of width) The length of the rectangle after it is increase by 2 inches will be l+2=3w+2 Width after it is increased by 1 inch will be w+1. Perimeter=(3w+2)+(w+1)+(3w+2)+(w+1) 62=8w+6 Solve this to find w. Then plug in that value of w in l=3w to find l.

OpenStudy (anonymous):

yes @chilimath

OpenStudy (anonymous):

@kaiteDID You don't always have to write out the words that @ajprincess included in her good explanation, but you should learn to think through the problem like that to help you set it up correctly... say those words in your mind as you set up the equation. Notice that the math all worked out the same as @chilimath showed, but if you aren't clear on how to get started, think about it like @ajprincess explained. Good luck!!